OpenStudy (anonymous):

Write the following in expanded form using logarithmic properties log5 x^3y ---- z^4

OpenStudy (anonymous):

can someone please help me

jimthompson5910 (jim_thompson5910):

The problem is \[\Large \log_{5} \left(\frac{x^3y}{z^4}\right)\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok we'll need to use these rules given here http://www.purplemath.com/modules/logrules.htm

jimthompson5910 (jim_thompson5910):

do those rules look familiar?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what's the first step? any ideas?

OpenStudy (anonymous):

would you do log5 X^3 +log5 Y - log5 Z^4

OpenStudy (anonymous):

it just looks wrong for some reason though

jimthompson5910 (jim_thompson5910):

you're very close to the final answer

OpenStudy (anonymous):

where did i go wrong on it? can you please explain

jimthompson5910 (jim_thompson5910):

Here's the full step by step picture (you'll see that you're on the right track)

jimthompson5910 (jim_thompson5910):

Use the log rules given on this page http://www.purplemath.com/modules/logrules.htm to get \[\Large \log_{5} \left(\frac{x^3y}{z^4}\right)\] \[\Large \log_{5} \left(x^3y\right)-\log_{5} \left(z^4\right) \ ... \ \text{Use rule 2}\] \[\Large \log_{5} \left(x^3\right)+\log_{5} \left(y\right)-\log_{5} \left(z^4\right) \ ... \ \text{Use rule 1}\] \[\Large 3\log_{5} \left(x\right)+\log_{5} \left(y\right)-4\log_{5} \left(z\right) \ ... \ \text{Use rule 3}\]

jimthompson5910 (jim_thompson5910):

So you just needed to use rule 3 to pull those exponents down

OpenStudy (anonymous):

ohh okay that makes sense thank you

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

one more quick question for this question log7p^6q^3r^1/4

OpenStudy (anonymous):

would it be like 6log7p + 3log7q + 1/4 log7r

jimthompson5910 (jim_thompson5910):

I'm assuming the original expression is \[\Large \log_{7} \left(p^{6}q^{3}r^{1/4}\right)\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Then \[\Large 6\log_{7} \left(p\right)+3\log_{7} \left(q\right)+\frac{1}{4}\log_{7} \left(r\right)\] is the correct expansion

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome